The ratio of men to women in elementary education courses is 2 to 7. If there are 270 students in an elementary education course, how many are women? What percent of the elementary education students are men?
Found 2 solutions by FrankM, lenny460:
Answer by FrankM(1040)

The ratio simply means that 2 of every nine are men. So 7 of 9 are women.
7/9 * 270 = 210 women.
2/9 = .22... (2 repeats)

Answer by lenny460(1073)

The ratio of men to women in elementary education courses is 2 to 7. If there are 270 students in an elementary education course, how many are women ? What percent of the elementary education students are men ?


7/9 * 270 = 7 * 30 = 210

Answer:
There are 210 women.

Percentage of men:
There are 60 men.
Therefore:
60/270 * 100 = 20/9 * 10 = 200/9 = 22.2%


Answer:
22.2%








Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com